If it takes 10minutes for the temperature of initial to fall from 100°c to 80°c within a surround temperature of 30°c. Calculate the Temperature of the initial after 30minutes


1
Expert's answer
2019-09-19T09:47:36-0400

We have:


"T-T_a=A\exp{(-kt)}"

The initial conditions:


"100-30=A\to A=70\degree C"

Now,

"80-30=70\exp{(-k(10))}"

"k=0.1\ln{1.4}"

So, temperature of the initial after 30 minutes:

"T=30+70\exp{(-0.1\ln{1.4}(30))}=55.5\degree C"